OpenStudy (anisah.21a):

log(21x^3)−log(7x^2)=3 how to solve for x?

jimthompson5910 (jim_thompson5910):

first you'll use the log rule log(A) - log(B) = log(A/B)

jimthompson5910 (jim_thompson5910):

using that rule turns `log(21x^3)−log(7x^2)` into what?

OpenStudy (anisah.21a):

log(21^3/7x^2)=3

OpenStudy (anisah.21a):

21/7=3 log10(3x)=3

jimthompson5910 (jim_thompson5910):

good by default, logs are in base 10. So you don't need to state the base if it's base 10

jimthompson5910 (jim_thompson5910):

log(3x) = 3

OpenStudy (anisah.21a):

10^3/3=3x/3 cancels out

OpenStudy (anisah.21a):

answer is x=10^3/3

jimthompson5910 (jim_thompson5910):

yep or 1000/3

OpenStudy (anisah.21a):

even though i just solved it myself lmao

jimthompson5910 (jim_thompson5910):

nice work